# Application of Large Language Models to Quantum State Simulation

### v1

@@ I Introduction
The structure of this paper is as follows: Section II provides a brief overview of the LLM framework used in this study and details the training process for simulating quantum circuits with LLMs. Section III presents numerical results and analyses, focusing on the single qubit quantum simulator model (Section III.1) and the two qubit quantum simulator model (Section III.2). Section IV discusses the expansion of the simulator, demonstrating how to construct a three qubit quantum circuit using a two qubit quantum simulator. Section V addresses the application of the quantum simulator under practical noise conditions and investigates its performance through experiments on actual quantum devices. Finally, Section VI summarizes the main conclusions of this study and explores potential directions for future research in this field.
@@ II Methodology
In this study, we propose a complex machine learning framework based on large language models (LLMs) for modeling and predicting the relationship between quantum circuit parameters and the corresponding quantum states. As an advanced pattern recognition tool, LLMs have a powerful ability to represent sequential data and can efficiently capture complex correlations between inputs and outputs through self-attention mechanisms. This feature makes LLMs particularly suitable for this study, as the mapping between quantum circuits and quantum states can be regarded as a complex sequence transformation problem. As shown in Fig. 1, the entire process involves using quantum circuit rotation gate parameters as input features, which are mapped by the LLM to the corresponding quantum state distribution. In this framework, the rotation angles of the quantum gates are extracted as continuous variables, and the combination of these variables determines the system’s final quantum state. During model training, we use these quantum circuit parameters and the target quantum states as training data and improve the prediction accuracy and stability by optimizing the loss function.
As shown in Fig. 2, the model structure proposed in this study includes encoder and decoder components[18, 19, 9]. The encoder is responsible for converting the input quantum circuit parameters into a concise fixed-length representation to capture the latent relationships in the input data. The decoder is used to generate the output target sequence, which can be either quantum state vectors or density matrices. The collaboration between the encoder and decoder enables the model to extract key information from the input features and effectively utilize it when generating target parameters.
Our model employs a self-attention mechanism[20], which is capable of dynamically adjusting weights to focus on the most relevant parts of the input data for the current prediction. Through the self-attention mechanism, the model can efficiently process the sequence of input quantum circuit parameters and generate the corresponding target quantum states. In our approach, the continuous parameters of quantum circuits, such as the rotation gate angles, are treated as input sequences. The model uses a multi-layer neural network to map these parameters into learnable representations, eventually generating accurate predictions of the target quantum states.
To efficiently train this framework, we rely on the design of the quantum circuit structure and the preparation of high-quality training data. The training data not only includes various combinations of quantum circuits but also covers the target quantum states or density matrices corresponding to these circuits, which are crucial for the model to learn the complexity of quantum state distributions. During training, we use a technique known as ”teacher forcing” to accelerate convergence. This technique is commonly used in sequence generation tasks, where the actual labels rather than the model’s predicted output from the previous timestep are used as input during training, thereby improving the model’s stability and convergence speed.
Furthermore, to enhance the model’s autoregressive capability in quantum state simulation tasks, we introduced an autoregressive mechanism in the decoder. Autoregression is a time series model that uses previous outputs as inputs to iteratively predict the next value in the sequence. In our model, this mechanism helps generate continuous quantum state predictions, allowing the model to better handle complex quantum systems.
The goal of the model is to minimize the mean squared error (MSE), a loss function that measures the difference between the model’s predictions and the actual quantum states. The formula for MSE is as follows:
@@
where $y_{i}$ represents the actual value, $\hat{y}_{i}$ represents the predicted value, and $n$ is the number of samples. By minimizing the mean squared error, we can improve the model’s prediction accuracy and generalization ability, ensuring that it can still provide high-quality quantum state predictions for previously unseen quantum circuits.
III Result
@@
Therefore, we use $\theta_{1},\theta_{2},\theta_{3}$ and $a_{00},a_{01},b_{01},a_{11}$ As the training set parameters for the 1-qubit noisy scenario. Due to the complexity of evolution in the presence of noise, we randomly generated more data and preprocessed it to train our model. The preprocessed data was then fed into the designed LLM for training.
@@
To extend the simulation to larger quantum systems, we constructed a 2-qubit quantum simulator that operates using 2 qubits. Figure 8 illustrates the schematic for obtaining special 2-qubit quantum circuits. Each qubit undergoes independent single-qubit rotation gates $U(\theta_{1},\theta_{2},\theta_{3})$ and $U(\theta_{4},\theta_{5},\theta_{6})$, followed by a CNOT gate to introduce quantum entanglement. Finally, the qubits are acted upon by single-qubit gates $U(\theta_{7},\theta_{8},\theta_{9})$ and $U(\theta_{10},\theta_{11},\theta_{12})$. This process can be represented as:
$U(\theta_{7},\theta_{8},\theta_{9})\otimes(U(\theta_{10},\theta_{11},\theta_{12})CNOTU(\theta_{1},\theta_{2},\theta_{3})$ (7)
$\otimes(U(\theta_{4},\theta_{5},\theta_{6})\begin{pmatrix}1\\ 0\end{pmatrix}\otimes\begin{pmatrix}1\\ 0\end{pmatrix}.$
The entire circuit generates an arbitrary two-qubit entangled state $|\varphi\rangle$ that conforms to our circuit structure. The quantum state vector of this entangled state is:
@@
IV Expansion of Quantum Simulator
In the previous section, we introduced a method to create a model for arbitrary 2-qubit circuits by combining two models. For 3-qubit circuits, we can extend this approach to simulate arbitrary 3-qubit quantum circuits using the same models. This involves using the 2-qubit model to simulate 2-qubit circuits and extending it to simulate 3-qubit circuits without needing additional training for a separate 3-qubit quantum circuit model.
@@ 3. Dashed Box 3: $q_{02}\otimes q_{1}$
V The model’s simulation under realistic noise conditions
@@
VI Conclusion
In this study, we introduce an innovative approach to quantum state simulation by leveraging the capabilities of LLMs. Our research has led to the development of a novel quantum simulator, rigorously tested through various experiments, including Grover’s algorithm and the Variational Quantum Eigensolver (VQE). The results consistently align with theoretical expectations, demonstrating the simulator’s reliability in both noise-free and noisy quantum circuit simulations.
A significant advancement in our work is the extension of the model from two-qubit systems to three-qubit quantum circuits, marking a pioneering exploration of higher-dimensional quantum systems. This extension highlights the simulator’s resource efficiency, enabling accurate replication of quantum computer experimental outcomes, encompassing quantum state vectors and density matrices, rather than merely predicting probability distributions.
Looking ahead, we anticipate that our LLM-based quantum simulator will play a pivotal role in the field of quantum computing. It offers the potential to simulate the outcomes of actual quantum circuits with reduced time and computational resource demands. Moreover, its inherent scalability from lower to higher dimensions significantly minimizes the costs associated with quantum simulations. This simulator presents itself as a valuable tool for pre-experimental testing, lowering experimental costs and technological barriers. Additionally, it holds promise as a powerful instrument for studying complex quantum systems, thereby accelerating advancements in quantum theory and applications.
VII Acknowledgement

### v2

@@ I Introduction
The structure of this paper is as follows: Section II provides a brief overview of the LLM framework used in this study and details the training process for simulating quantum circuits with LLMs. Section III presents the numerical results and analysis, focusing on the single-qubit quantum simulator model (Section III.1) and the two-qubit quantum simulator model (Section III.2). Section III.3 discusses the extension of the simulator, demonstrating how to construct a three-qubit quantum circuit using the two-qubit quantum simulator and presenting experimental results that test this extension method. Section III.4 addresses the application of the quantum simulator under realistic noise conditions, investigating its performance through experiments conducted on actual quantum devices. Finally, Section IV summarizes the limitations of this study and discusses potential directions for future research.
@@ II Methodology
In this study, we propose and validate a framework based on large language models (LLMs) for modeling and predicting the complex relationships between quantum circuit parameters and their corresponding quantum states. As advanced natural language processing tools, LLMs are widely used in fields such as natural language generation, translation, and dialogue due to their powerful capabilities in processing sequential data. By treating the rotation gate parameters (e.g., angles) of quantum circuits as input sequences and utilizing the self-attention mechanism of LLMs to model their interdependencies, we successfully applied this framework to the prediction of quantum states.
Specifically, the input to the model consists of the rotation angles of each quantum gate, while the output from the LLM includes the predicted quantum state, represented by the quantum state vector and the distribution of the density matrix. As illustrated in Figure 1, this figure visually depicts the complete process from the input of quantum circuit parameters to the LLM’s output of the corresponding quantum state. The core of this process lies in the LLM’s ability to efficiently capture the complex interdependencies of quantum circuit parameters through its multi-layer self-attention mechanism, thereby achieving accurate modeling of quantum states in a high-dimensional space.
@@
Compared to traditional machine learning methods, such as neural networks or classical regression models, LLMs exhibit stronger adaptability and predictive capability in quantum circuit simulations. While neural networks can handle certain nonlinear relationships, they often face limitations in generalization or computational efficiency when dealing with high-dimensional, complex quantum systems. Classical regression models, on the other hand, are typically used for linear problems and struggle to capture the nonlinear and complex interactions within quantum circuits. In contrast, LLMs can globally capture the interactions between quantum gates through their self-attention mechanism, and with their deep, stacked structure, they demonstrate superior predictive and generalization capabilities.
As shown in Figure 2, the proposed model structure includes two components: an encoder and a decoder [18, 19, 9]. The encoder’s role is to convert the input quantum circuit parameters into concise fixed-length representations, capturing the latent relationships within the input data. The decoder is used to generate the output target sequence, which can be a quantum state vector or a density matrix. The collaboration between the encoder and decoder allows the model to extract key information from the input features and effectively utilize this information when generating target parameters, thereby improving the accuracy and reliability of the generated quantum states.
@@
Our model employs a self-attention mechanism [20], which dynamically adjusts weights to focus on the most relevant parts of the input data needed for the current prediction. Through the self-attention mechanism, the model can efficiently process the input sequence of quantum circuit parameters and generate the corresponding target quantum states. In the specific implementation, the continuous parameters of the quantum circuit (such as rotation angles) are treated as an input sequence, with the model using a multi-layer neural network to map these parameters into learnable representations, ultimately producing accurate predictions of the target quantum states.
To train this framework effectively, we rely on the reasonable design of the quantum circuit structure and the preparation of high-quality training data. This training data not only includes combinations of various quantum circuits but also covers the corresponding target quantum states or density matrices, which are crucial for the model to learn the complexities of quantum state distributions. During the training process, we employed a technique known as ”teacher forcing” to accelerate convergence and improve model stability. This technique is commonly used in sequence generation tasks, where actual labels, rather than the model’s predictions from the previous time step, are used as input during training, effectively enhancing the training speed and accuracy of the model.
Additionally, to enhance the model’s autoregressive capabilities in quantum state simulation tasks, we introduced an autoregressive mechanism in the decoder. Autoregression is a time series model that uses previous outputs as inputs to iteratively predict the next value in the sequence. In our model, this mechanism not only improves the model’s ability to generate continuous quantum states but also enhances its performance when handling complex quantum systems.
The objective of the model is to minimize the mean squared error (MSE), which serves as a loss function measuring the difference between the model’s predicted values and the actual quantum states. The formula for MSE is as follows:
@@
where $y_{i}$ represents the actual value, $\hat{y}_{i}$represents the predicted value, and $n$ is the number of samples. By minimizing the mean squared error, the model’s prediction accuracy and generalization capability are improved, ensuring that it can provide high-quality quantum state predictions for previously unseen quantum circuits.
III Results
@@
Therefore, we select $\theta_{1},\theta_{2},\theta_{3}$ , as well as $a_{00},a_{01},b_{01},a_{11}$ as the training parameters in the 1-qubit noisy scenario. Due to the complexity of evolution in the presence of noise, we randomly generated more data and preprocessed it to train our model. The preprocessed data was then fed into the designed LLM for training.
@@
To extend the simulation to larger quantum systems, we constructed a 2-qubit quantum simulator that operates using 2 qubits. Figure 8 illustrates the schematic for obtaining special 2-qubit quantum circuits. Each qubit undergoes independent single-qubit rotation gates $U(\theta_{1},\theta_{2},\theta_{3})$ and $U(\theta_{4},\theta_{5},\theta_{6})$, followed by a CNOT gate to introduce quantum entanglement. Finally, the qubits are acted upon by single-qubit gates $U(\theta_{7},\theta_{8},\theta_{9})$ and $U(\theta_{10},\theta_{11},\theta_{12})$. This quantum circuit generates an arbitrary two-qubit entangled state $|\varphi\rangle$ that conforms to our circuit structure.
The quantum state vector of this entangled state is:
@@
III.3 Expansion of Quantum Simulator
In the previous section, we introduced a method for constructing a simulator for arbitrary two-qubit circuits by combining two models. Building on this framework, we further explored how to extend this method to simulate three-qubit circuits without incurring additional training costs. By leveraging the existing two-qubit models, we can effectively construct a three-qubit simulator without the need to retrain a separate three-qubit model. This extension not only reduces training time and computational resource consumption but also significantly enhances the model’s scalability, making it applicable to more complex quantum circuit simulations.
@@ 3. Dashed Box 3: $q_{02}\otimes q_{1}$
Figure 17: This figure shows the fidelity between the model results and the actual quantum circuit results for different samples. The horizontal axis ”Sample Index” represents the index of each test sample, where each point corresponds to a different random 3-qubit circuit and its related quantum circuit result. The vertical axis ”Fidelity Between Model and Circuit” represents the fidelity between the density matrix predicted by the model and the density matrix output by the actual quantum circuit. As can be observed from the figure, the fidelity for each test sample is above 93%, indicating the accuracy of the model in simulating quantum states.
To validate the performance of the model, we designed and conducted a series of experiments simulating several randomly generated three-qubit circuits. We compared the results generated by the model with the outcomes from actually running the same quantum circuits. The core metric in these experiments was the fidelity between the density matrices, which measures the similarity between the quantum states generated by the model and those produced by real hardware. Figure 17 presents these fidelity results, where the horizontal axis represents different random three-qubit circuit samples, and the vertical axis indicates the fidelity between the model-generated density matrices and the actual results. The experimental results show that the model exhibits high fidelity across all test samples, with fidelities consistently exceeding 93%, demonstrating the model’s high accuracy and stability in quantum state simulation. Regardless of the complexity of the circuit structure, the model exhibited highly consistent performance, further confirming its adaptability to different quantum state evolution pathways.
[…]
Figure 18: Parameterized circuit for a three-qubit Variational Quantum Eigensolver (VQE). Multiple rotation gates $Rx(\theta_{i})$ and $Rz(\theta_{j})$ are applied to each qubit line, with adjustable parameters $\theta$ to optimize the quantum state. The circuit also includes controlled-NOT (CNOT) gates to introduce entanglement between qubits. This structure allows the circuit to minimize the expectation value of the Hamiltonian by tuning the parameters, thereby approximating the ground state energy of the system.
To further assess the capabilities of this extended model, we conducted an experiment based on the Variational Quantum Eigensolver (VQE) to evaluate the model’s performance in solving the ground-state energy problem of complex quantum systems. The VQE experiment aims to approximate the ground-state energy of a target Hamiltonian by optimizing a parameterized quantum circuit. The VQE circuit, as shown in Figure 18, includes several rotation gates $Rx(\theta_{i})$ and $Rz(\theta_{j})$ and introduces quantum entanglement via controlled-NOT (CNOT) gates, ensuring that the circuit captures the complex quantum state evolution. The target Hamiltonian used in the experiment is:
$H=X_{1}\otimes X_{2}\otimes X_{3}+Y_{1}\otimes Y_{2}\otimes Y_{3}+Z_{1}\otimes Z_{2}\otimes Z_{3},$ (16)
where $X_{i}$, $Y_{i}$, and $Z_{i}$ are Pauli operators acting on the $i$-th qubit. Theoretically, the ground-state energy of this Hamiltonian is $-1.4$. As shown in Figure 19, after VQE optimization, the model estimated the ground-state energy to be $-1.33$. Although there is a deviation from the theoretical value, the result still demonstrates the model’s ability to approximate complex quantum systems, proving its effectiveness in variational quantum algorithms.
Figure 19: The figure illustrates the energy convergence process in the Variational Quantum Eigensolver (VQE) experiment. The horizontal axis represents the number of optimization iterations, while the vertical axis shows the estimated ground state energy of the system after each iteration. As the number of iterations increases, the parameters are gradually optimized, and the energy approaches the ground state energy of the target Hamiltonian.
Overall, the extended three-qubit simulator, based on the two-qubit model, exhibits superior performance in multiple aspects. First, by extending the two-qubit architecture, the model efficiently simulates three-qubit circuits without retraining, significantly improving simulation efficiency. Second, the experimental results show that the model achieves high accuracy and stability in both quantum state simulation and variational quantum eigensolver tasks, further validating its potential in simulating more complex quantum systems. This lays a solid foundation for future applications of large language models in simulating complex quantum circuits and provides new directions for research in quantum computing.
III.4 The model’s simulation under realistic noise conditions
@@
IV Discussion
This study integrates large language models (LLMs) to develop a novel quantum state simulator, which was extensively tested using Grover’s algorithm and the Variational Quantum Eigensolver (VQE). The test results show that the simulator exhibits high reliability when handling quantum circuits in both noise-free and noisy environments, aligning well with theoretical expectations.
Despite these significant achievements, the simulator still has some limitations. Firstly, although it performed well in three-qubit experiments, when scaling to larger numbers of qubits, the model may encounter resource bottlenecks, and efficiency optimization is required. Secondly, while the model maintains high accuracy under current noise conditions, its precision may be affected when dealing with more complex noise models. Therefore, future research should focus on enhancing the model’s robustness to address a wider variety of quantum noise environments.
The study also demonstrates that the LLM-based simulator can efficiently reproduce both quantum state vectors and density matrices, enabling comprehensive simulations of complex quantum systems. This capability is particularly valuable for quantum computing experimental simulations, as it significantly reduces resource consumption during the initial stages of experiments. By performing efficient simulations in advance, researchers can better assess the performance of quantum circuits in practical scenarios, optimize experimental design, and minimize trial-and-error costs.
The introduction of LLMs offers new research perspectives for the quantum computing field, particularly in resource-constrained situations. This study highlights the tremendous potential of LLMs in quantum state simulation, especially in terms of model scalability. Additionally, LLMs provide researchers with an alternative pre-experimental testing method, lowering the cost barriers of expensive quantum computing experiments. As the model continues to expand, it is expected to more easily predict the complexity of actual quantum computing experiments. More importantly, the high computational speed of LLMs opens new possibilities for improving and optimizing quantum algorithms, particularly those involving large numbers of quantum gate operations.
Future research could focus on several areas: first, optimizing the structure of LLMs to ensure their continued efficiency when simulating larger numbers of qubits; second, further improving the simulator’s robustness in complex quantum noise environments, particularly by fine-tuning the model for more detailed simulations of noise conditions in real quantum devices. Furthermore, given the flexibility of LLMs, future studies could explore their application in a broader range of quantum algorithms, such as simulating quantum many-body systems and quantum machine learning. These research directions would not only enhance the multifunctionality of quantum simulators but also drive widespread applications in both theoretical and practical aspects of quantum computing.
V Acknowledgement